# A Vision for Trustworthy, Fair, and Efficient Socio-Technical Control using Karma Economies

### v1

@@ 4.4 Model refinement with Dynamic Population Games (DPGs)
The KE $(\bm{\chi^{\star}},\bm{b^{\star}})$ provides a high-level description of the karma economy. From $(\bm{\chi^{\star}},\bm{b^{\star}})$, one can construct explicit bidding behaviors for each each time-step $t\in{\mathbb{N}}$, in which user $i\in{\mathcal{N}}$ in urgency $u_{t,i}\in{\mathcal{U}}_{i}$ bids $b_{t,i,j}=b^{\star}_{j}$ (respectively, $b_{t,i,j}=0$) on resources $j\in{\mathcal{M}}$ with probability $\chi^{\star}_{i,j}(u_{i})$ (respectively, $1-\chi^{\star}_{i,j}(u_{i})$). However, since the KE respects the karma budget in a long-run expected average sense, a user may not always have enough karma to realize its bid ex-post. Moreover, the KE assumes that users are infinitely far-sighted while users may prioritize immediate rewards over future ones. For this purpose, a complementary Dynamic Population Game (DPG) model was developed that gives fine grained insights on the karma economy (Elokda et al., 2024b, c). DPGs are a tractable subclass of mean-field games that enable computing complex and high-dimensional Stationary Nash Equilibria (SNE) (Elokda et al., 2024a). The SNE is a finer-grained solution concept than the KE that fully accounts for the karma budget in every time-step, considers discounted future rewards, and provides detailed bidding policies mapping both urgency and karma to (potentially probabilistic) bids. Moreover, in addition to optimal bidding policies, the SNE predicts the stationary distribution of karma in the population, which holds more information than the stationary bids $\bm{b^{\star}}$. For example, the SNE allows to quantify the fraction of times users are expected to spend below a certain karma level, as well as whether ‘karma wealth inequalities’ can arise (Elokda et al., 2024b). The DPG model was applied in multiple extensions, including endogenous redistribution shares that depend on the users’ actions, e.g., at what time they depart during the morning rush hour (Elokda et al., 2024c), elastic population of users (Elokda et al., 2023), and pairwise or otherwise limited resource contests such as in autonomous driving (Elokda et al., 2024b; Chavoshi et al., 2024). However, the DPG model is computational in nature and less amenable to analysis, for which the high-level KE is advantageous.
@@ 4.5 Related mechanisms
The Competitive Equilibrium (CE) is a cornerstone of economic theory that models how users trade resources in large markets (Arrow and Debreu, 1954). In full generality, the budget of each user is its initial endowment of resources, or the resources it contributes to the market, and the CE is a set of resource prices at which each user is best off selling part or all of its initial endowment and buying resources from others, effectively facilitating efficient trade. Fisher markets are a more closely related special case, in which the resources are public or owned by a central entity and each user is endowed with a budget of artificial currency, representing its access rights to resources (Vazirani, 2007). In the case of linear and separately additive utilities, Fisher market CE maximize Nash welfare (Vazirani, 2007, Theorem 5.1). Moreover, when the resources are indivisible, similar schemes go by the name of pseudomarkets (Hylland and Zeckhauser, 1979) and have been famously applied in the context of allocating course seats to students in business schools (Budish, 2011).
These mechanisms that budget over space rely on there being a large variety of resources and heterogeneity in needs to attain efficiency gains. If the users are homogeneous, or if there is a single (most desirable) resource, these mechanisms yield trivial equal splits of resources (if all users have equal access rights). The time dimension is thus vital to realize efficiency gains beyond those attainable with space alone.
@@
One may argue that time is not fundamentally different from space: if a resource is repeatedly allocated over $T$ time-steps, this can be ‘converted’ to space by allocating $T$-copies of the resource in one shot. But the time dimension deserves a separate treatment for at least two reasons. First is the online nature of time: users must decide their allocation one time-step at a time, with imperfect knowledge of their future preferences. Second, the analogy breaks down if the resources are allocated indefinitely, i.e., $T\rightarrow\infty$, as this would necessitate infinite initial endowments in the classical spacial models.
Works on budgeting over time draw their roots in the linking mechanism by Jackson and Sonnenschein (2007) discussed in Section 2.3. The linking mechanism was extended to Markovian rather than i.i.d. preference dynamics in Escobar and Toikka (2013), which also assumes public knowledge of the dynamic processes and tailors the mechanism to these processes over finite time periods. The most closely related mechanisms that do not require public knowledge of the preference dynamics are those based on artificial currency, also known as online Fisher markets as they form the natural extension of static Fisher markets (Gorokh et al., 2021; Gao et al., 2021; Banerjee et al., 2023; Jalota and Ye, 2025; Lin et al., 2025). In these mechanisms, each user receives an initial endowment of artificial currency to spend over a finite horizon. While research on online Fisher markets remains active, we remark that they share the same shortcoming of linking mechanisms of being ‘open-loop.’ The lack of feedback, as facilitated in karma economies by the redistribution of karma, makes them not well-suited for infinite resource repetitions; and sensitive to transient shocks or bidding mistakes that could prematurely deplete the budget with no opportunity for recovery. Figure 8 illustrates this distinguishing feature of karma.
{mdframed}
@@ 5 Outlook: To Couple or Not to Couple? A Smart City Run on Multi-Karma Economies
(Coupling Karma Economies Benefits All) Suppose that there are two resources ${\mathcal{M}}={1,2}$, and let $\bar{r}_{i}^{\star,1}$ (respectively, $\bar{r}_{i}^{\star,2}$) denote user $i$’s long-run expected average reward derived from resource $1$ (respectively, resource $2$) at the KE of two separate karma economies. If a KE of the combined karma economy exists, then it satisfies, for all $i\in{\mathcal{N}}$
$\bar{r}_{i}^{\star}\geq\bar{r}_{i}^{\star,1}+\bar{r}_{i}^{\star,2},$
where $\bar{r}_{i}^{\star}$ denotes user $i$’s long-run expected average reward derived from both resources in the combined economy.
@@ 6 Conclusion
Given the frequency and complexity of karma bidding, it would be unreasonable to expect citizens of the future smart city to participate in karma economies directly. The design of simple user interfaces, as well as automated bidders that bid on the users’ behalf in real-time, are therefore promising directions for future research. In this direction, preliminary research has subjected karma economies to the behavioral test with real humans (Elokda et al., 2024e), as well as investigated adopting automated bidders for monetary ad auctions to the non-monetary karma setting (Berriaud et al., 2024). The active research area of multi-agent reinforcement learning (Zhang et al., 2021) is also relevant for this challenge.
@@ A.2 Proof of Proposition 4
[…]
@@
Finally, Proposition 4.7 is immediate from the budget balance condition of KE when there is a single resource =${\mathcal{M}}=\{1\}$.
@@ A.3 Proof of Theorem 5
[…]

### v2

@@ 4.4 Model refinement with Dynamic Population Games (DPGs)
The KE $(\bm{\chi^{\star}},\bm{b^{\star}})$ provides a high-level description of the karma economy. From $(\bm{\chi^{\star}},\bm{b^{\star}})$, one can construct explicit bidding behaviors for each time-step $t\in{\mathbb{N}}$, in which user $i\in{\mathcal{N}}$ in urgency $u_{t,i}\in{\mathcal{U}}_{i}$ bids $b_{t,i,j}=b^{\star}_{j}$ (respectively, $b_{t,i,j}=0$) on resources $j\in{\mathcal{M}}$ with probability $\chi^{\star}_{i,j}(u_{i})$ (respectively, $1-\chi^{\star}_{i,j}(u_{i})$). However, since the KE respects the karma budget in a long-run expected average sense, a user may not always have enough karma to realize its bid ex-post. Moreover, the KE assumes that users are infinitely far-sighted while users may prioritize immediate rewards over future ones. For this purpose, a complementary Dynamic Population Game (DPG) model was developed that gives fine grained insights on the karma economy (Elokda et al., 2024b, c). DPGs are a tractable subclass of mean-field games that enable computing complex and high-dimensional Stationary Nash Equilibria (SNE) (Elokda et al., 2024a). The SNE is a finer-grained solution concept than the KE that fully accounts for the karma budget in every time-step, considers discounted future rewards, and provides detailed bidding policies mapping both urgency and karma to (potentially probabilistic) bids. Moreover, in addition to optimal bidding policies, the SNE predicts the stationary distribution of karma in the population, which holds more information than the stationary bids $\bm{b^{\star}}$. For example, the SNE allows to quantify the fraction of times users are expected to spend below a certain karma level, as well as whether ‘karma wealth inequalities’ can arise (Elokda et al., 2024b). The DPG model was applied in multiple extensions, including endogenous redistribution shares that depend on the users’ actions, e.g., at what time they depart during the morning rush hour (Elokda et al., 2024c), elastic population of users (Elokda et al., 2023), and pairwise or otherwise limited resource contests such as in autonomous driving (Elokda et al., 2024b; Chavoshi et al., 2024). However, the DPG model is computational in nature and less amenable to analysis, for which the high-level KE is advantageous.
@@ 4.5 Related mechanisms
The Competitive Equilibrium (CE) is a cornerstone of economic theory that models how users trade resources in large markets (Arrow and Debreu, 1954). In full generality, the budget of each user is its initial endowment of resources, or the resources it contributes to the market, and the CE is a set of resource prices at which each user is best off selling part or all of its initial endowment and buying resources from others, effectively facilitating efficient trade. Fisher markets are a more closely related special case, in which the resources are public or owned by a central entity and each user is endowed with a budget of artificial currency, representing its access rights to resources (Vazirani, 2007). In the case of linear and separately additive utilities, Fisher market CE maximize Nash welfare (Vazirani, 2007, Theorem 5.1). Moreover, when the resources are indivisible, similar schemes go by the name of pseudomarkets (Hylland and Zeckhauser, 1979) and have been famously applied in the context of allocating course seats to students in business schools (Budish, 2011). These mechanisms that budget over space rely on there being a large variety of resources and heterogeneity in needs to attain efficiency gains. If the users are homogeneous, or if there is a single (most desirable) resource, these mechanisms yield trivial equal splits of resources (if all users have equal access rights). The time dimension is thus vital to realize efficiency gains beyond those attainable with space alone.
@@
One may argue that time is not fundamentally different from space: if a resource is repeatedly allocated over $T$ time-steps, this can be ‘converted’ to space by allocating $T$-copies of the resource in one shot. But the time dimension deserves a separate treatment for at least two reasons. First is the online nature of time: users must decide their allocation one time-step at a time, with imperfect knowledge of their future preferences. Second, the analogy breaks down if the resources are allocated indefinitely, i.e., $T\rightarrow\infty$, as this would necessitate infinite initial endowments in the classical spacial models. Works on budgeting over time draw their roots in the linking mechanism by Jackson and Sonnenschein (2007) discussed in Section 2.3. The linking mechanism was extended to Markovian rather than i.i.d. preference dynamics in Escobar and Toikka (2013), which also assumes public knowledge of the dynamic processes and tailors the mechanism to these processes over finite time periods. The most closely related mechanisms that do not require public knowledge of the preference dynamics are those based on artificial currency, also known as online Fisher markets as they form the natural extension of static Fisher markets (Gorokh et al., 2021; Gao et al., 2021; Banerjee et al., 2023; Jalota and Ye, 2025; Lin et al., 2025). In these mechanisms, each user receives an initial endowment of artificial currency to spend over a finite horizon. While research on online Fisher markets remains active, we remark that they share the same shortcoming of linking mechanisms of being ‘open-loop.’ The lack of feedback, as facilitated in karma economies by the redistribution of karma, makes them not well-suited for infinite resource repetitions; and sensitive to transient shocks or bidding mistakes that could prematurely deplete the budget with no opportunity for recovery. Figure 8 illustrates this distinguishing feature of karma.
@@ 5 Outlook: To Couple or Not to Couple? A Smart City Run on Multi-Karma Economies
(Coupling Karma Economies Benefits All) Suppose that there are two resources ${\mathcal{M}}={1,2}$, and let $\bar{r}_{i}^{\star,1}$ (respectively, $\bar{r}_{i}^{\star,2}$) denote user $i$’s long-run expected average reward derived from resource $1$ (respectively, resource $2$) at the KE of two separate karma economies. If a KE of the combined karma economy exists, then it satisfies, for all $i\in{\mathcal{N}}$ $\bar{r}_{i}^{\star}\geq\bar{r}_{i}^{\star,1}+\bar{r}_{i}^{\star,2},$ where $\bar{r}_{i}^{\star}$ denotes user $i$’s long-run expected average reward derived from both resources in the combined economy.
@@ 6 Conclusion
Applications
As a novel general resource allocation method, and as pointed out in Section 5, karma economies could impact many socio-technical application domains, including the canonical examples of transportation, energy, bandwidth, data, etc.. It is particularly important to study how karma can contribute to climate actions and the pressing clean energy transition. There is a plethora of problems to explore, ranging from karma-based EV smart charging / peak-shaving; resource sharing in energy communities; fair ways to elicit demand response and provide flexibility for renewable generation; all the way to a completely re-imagined, karma-based future energy market with 100% renewable and publicly available clean energy.
@@
Given the frequency and complexity of karma bidding, it would be unreasonable to expect citizens of the future smart city to participate in karma economies directly. The design of simple user interfaces, as well as automated bidders that bid on the users’ behalf in real-time, are therefore important directions for future research that will enable scaling the present approach to more complex real-world settings. In this direction, preliminary research has subjected karma economies to the behavioral test with real humans (Elokda et al., 2024e), as well as investigated adopting automated bidders for monetary ad auctions to the non-monetary karma setting (Berriaud et al., 2024). The active research area of multi-agent reinforcement learning (Zhang et al., 2021) is also relevant for this challenge.
Non-stationary environments
A standing assumption of this paper is that resource capacities and reward structures are time-invariant, which allows studying the compact equilibrium concept of Karma Equilibrium (KE) expressed in terms of stationary long-run allocations and clearing bids. An important future extension thus regards relaxing this assumption and allowing resource capacities and reward structures to vary over time, e.g., due to weather conditions, seasonal changes, or disturbance events. A promising avenue in this direction, that was explored in Elokda et al. (2024d), is to introduce a global environment state characterizing changing conditions, and define KE in terms of stationary long-run allocations and clearing bids conditioned on the environment state.
@@ A.2 Proof of Proposition 4
$\Leftrightarrow\quad$ $\kappa^{\star}_{i}=\frac{\sum_{i^{\prime}\in{\mathcal{N}}}w_{i^{\prime}}}{w_{i}\sum_{j\in{\mathcal{M}}}b^{\star}_{j}\,c_{j}}\left(\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{\star}-\sum_{u_{i}\in{\mathcal{U}}_{i}}\sum_{j\in{\mathcal{M}}}\eta^{\star}_{i,j}(u_{i})\right).$
@@
Finally, Proposition 4.7 is immediate from the budget balance condition of KE when there is a single resource ${\mathcal{M}}=\{1\}$.
@@ A.3 Proof of Theorem 5
$\Leftrightarrow\quad$ $1-C=\frac{\sum_{j\in{\mathcal{M}}}\lambda^{*}_{j}\,c_{j}}{\sum_{i\in{\mathcal{U}}_{i}}w_{i}}.$